A wheel rotating at 490 rev/min accelerate uniformly at 0.5 rad/s2 for 30 second.


1 .find the angular velocity after 30s


2 . pulley peripheral speed


3 . angle turned through by a wheel during this time

1
Expert's answer
2022-05-15T17:02:26-0400

Given:

"\omega_0=2\pi*490*1/60=51.3\rm\: rad/s"

"\alpha=0.5\:\rm rad/s^2"

"t=30\:\rm s"


1.

"\omega_f=\omega_0+\alpha t=51.3+0.5*30=66.3\:\rm rad/s"

2. Not Enough Data


3.

"\theta=\omega_0t+\alpha t^2/2\\
=51.3*30+0.5*30^2/2=1764\:\rm rad"